OpenStudy (anonymous):

Write the expression as either the sine, cosine, or tangent of a single angle. sin ( pi / 2 ) * cos( pi / 7 ) + cos ( pi / 2 ) * sin (pi/ 7)

OpenStudy (psymon):

This is the result of a sum of sines formula. This is what you get when you have sin(a + b). So just take what you have and put it backwards into the sum of sines sin(a + b)

OpenStudy (anonymous):

So add pi over 2 and pi over 7 together

OpenStudy (psymon):

Pretty much xD

OpenStudy (anonymous):

Alright but first would I cross multiply them

OpenStudy (psymon):

Nope, it's just \[\sin(\frac{ \pi }{ 2 }+\frac{ \pi }{ 7 })\]

OpenStudy (anonymous):

Ok so it would be \[\sin ( \frac{ \pi }{ 9 } )\]

OpenStudy (psymon):

Well, we need to add the fractions like normal, common denominators and things like that.

OpenStudy (anonymous):

Alright would it be \[\sin ( 2 \pi / 9 ) \] then

OpenStudy (psymon):

Hmm \[\frac{ \pi }{ 2 }*\frac{ 7 }{ 7 } + \frac{ \pi }{ 7 }*\frac{ 2 }{ 2 }\] \[\frac{ 7\pi }{ 14 }+ \frac{ 2\pi }{ 14 } = \frac{ 9\pi }{ 14 }\] Kinda see how?

OpenStudy (anonymous):

yeah and this would be a sin expression

OpenStudy (psymon):

Yeah, itd be sin(9pi/14) and that would be your answer.

OpenStudy (anonymous):

Ok thanks

OpenStudy (psymon):

Mhm.